# On the Impact of Dataset Properties on Membership Privacy of Deep Learning

### v2

@@ 3.2 Experimental Results
Using the setting described above, we study how the number of classes and the number of shots affect the vulnerability (tpr at fpr as described in Section 2). We make the following observations:
  • A larger number of $S$ (shots) decrease the vulnerability in a power law relation as demonstrated in Figure 1(a). We provide further evidence of this in the Appendix (Figures A.1, A3 and A4).
  • Contrary, a larger number of $C$ (classes) increases the vulnerability as demonstrated in Figure 1(b) with further evidence in Figures A.2, A5 and A6 in the Appendix. However, the trend w.r.t. $C$ is not as clear as with $S$.
@@
Figure 1: MIA vulnerability (tpr at $\textsc{fpr}=0.001$) as a function of data set properities when attacking a ViT-B Head fine-tuned without DP on different data sets. The solid line displays the median and the error bars the minimum of the lower bounds and maximum of the upper bounds for the Clopper-Pearson CIs over multiple seeds (six for Figure 1(a) and 12 for Figure 1(b))
@@ 3.3 Model to Predict Data Set Vulnerability
[…]
@@
We utisee MIA results of ViT-B (see Table A3) as the training data. Based on the $R^{2}$ (coefficient of determination) score ($R^{2}=0.930$ for the model trained on $\textsc{fpr}=0.001$ data), our model fits the data extremely well. We provide further evidence for other fpr in Figure A.3 and Table A8 in the Appendix. Figure 2 shows the parameters of the prediction model fitted to the training data.
Figure 2: Coefficient values for different fpr when fitting a regression model based on Equation 2 fitted on data from ViT-B (Head) (Table A3). The error bars display the 95% confidence intervals based on Student’s t-distribution.
@@
  • R-50 (FiLM): Figure 3(b) shows that the prediction quality is good for R-50 (FiLM) models. These models are fine-tuned with parameter-efficient FiLM (Perez et al., 2018) layers (See Section A.1). Tobaben et al. (2023) demonstrated that FiLM layers are a competitive alternative to training all parameters. We supplement the MIA results of Tobaben et al. (2023) with own FiLM training runs. Refer to Table A7 in the Appendix.
@@
(b) Evaluating prediction performance for target models 1 left: fine-tuned with FiLM (data partially from Tobaben et al. (2023), see Table A7) 2 right: trained from scratch (data from Carlini et al. (2022))
@@
4 Understanding the MIA Prediction Model
In this section, we focus on understanding why the previously observed trends exist. Our goal is to formally prove a similar relationship as in the predictive model of Section 3.3.
4.1 A Simplified Model of Membership Inference
In this section, we construct a simplified model of membership inference that streamlines the data generation and shadow model training. We observe that the trends of the vulnerability of the simplified model is a good approximation of the trends previously observed with the image classification example.
Instead of training linear classifiers to classify images, we sample vectors on a high dimensional unit sphere and classify based on inner product with estimated class mean. This model is easier to analyse theoretically, and allows us to simulate arbitrary combinations of $S$ and $C$ without relying on the existence of a specific image classification downstream data set as especially high $C$ image classification downstream data sets that are computationally feasible to fine-tune on have small $S$. We generate the data and form the classifiers (which are our target models) as follows:
  1. 1.
    For each class, we first sample a true class mean $m_{c}$ on a high dimensional unit sphere that is orthogonal to all other true class means ($\forall i,j\in C:m_{i}\perp m_{j}\lor i=j$).
  2. 2.
  3. 3.
    For each "target model" we randomly choose a subset of size $CS$ from all generated vectors and compute per-class means $r_{c}$.
  4. 4.
    The computed mean is used to classify sample $(x,y)$ by computing the inner product $\langle x,r_{c}\rangle$ as a metric of similarity.
The attacker has to infer which vectors have been used for training the classifier. Instead of utilising the logits (like in the image classification example), the attacker can use the inner products of a point with the cluster means. In Figure 4 we compare the attacker’s performance as a function of $C$ and $S$ to the the previous image classification example.
This simplified model resembles the linear (Head) classifier. In the linear classifier, we find a matrix $A$ and biases $b$, to optimize the cross-entropy between the labels and logits $Az+b$, where $z$ denotes the feature space representation of the data. In the simplified model, the rows of $A$ are replaced by the cluster means and we do not include the bias term in the classification.
(a) as a function of $S$ (shots)
(b) as a function of $C$ (classes)
[…]
4.2 Theoretical Analysis
Similarly to the setting in Section 4.1, we focus on a classification model that aims to classify the sample $(x,y)$ based on the inner product $\langle x,r_{c}\rangle$, where $r_{c}$ is the mean of the training data points $x_{m}$ with $y_{m}=c$. However, we only focus on the scores of the target class, i.e. $c=y$, and therefore the score used of LiRA simplifies to $\langle x,r_{y}\rangle$.
The fpr for LiRA is given as
$\textsc{fpr}=\Pr\left(\frac{\mathcal{N}(t_{x};\mu_{\textrm{in}},\sigma_{\textrm{in}}^{2})}{\mathcal{N}(t_{x};\mu_{\textrm{out}},\sigma_{\textrm{out}}^{2})}>\exp(\tau)\mid x\notin\mathcal{D}_{\textrm{target}}\right),$ (3)
where $t_{x}=\langle x,r_{y}^{\textrm{(target)}}\rangle$ is the score for the target model, $\tau$ is the decision threshold for the classifier, and $\mu_{\textrm{in}},\mu_{\textrm{out}},\sigma_{\textrm{in}}$ and $\sigma_{\textrm{out}}$ are the LiRA statistics computed from the shadow models. We assume that $\sigma_{\textrm{in}}=\sigma_{\textrm{out}}=\sigma$, which allows to write Equation 3 as
fpr $=\Pr\left(\log\frac{\mathcal{N}(t_{x};\mu_{\textrm{in}},\sigma^{2})}{\mathcal{N}(t_{x};\mu_{\textrm{out}},\sigma^{2})}>\tau\mid x\notin\mathcal{D}_{\textrm{target}}\right)$ (4)
$=\Pr\left(-\frac{1}{2\sigma^{2}}(\mu_{\textrm{in}}-\mu_{\textrm{out}})\left[-2t_{x}+(\mu_{\textrm{in}}+\mu_{\textrm{out}})\right]>\tau\mid x\notin\mathcal{D}_{\textrm{target}}\right).$ (5)
From now on, we will denote the $-\frac{1}{2\sigma^{2}}(\mu_{\textrm{in}}-\mu_{\textrm{out}})\left[-2t_{x}+(\mu_{\textrm{in}}+\mu_{\textrm{out}})\right]$ with $\lambda_{\textsc{fpr}}$ when $x\notin\mathcal{D}_{\textrm{target}}$ and with $\lambda_{\textsc{tpr}}$ otherwise. Furthermore we will drop the condition $x\in\mathcal{D}_{\textrm{target}}$ or $x\notin\mathcal{D}_{\textrm{target}}$ from the fpr and tpr expressions
fpr $:=\Pr(\lambda_{\textsc{fpr}}>\tau)$ (6)
tpr $:=\Pr(\lambda_{\textsc{tpr}}>\tau).$ (7)
Lemma 1.
In the limit of infinite shadow models and with large enough $S$, the LiRA test statistics $\lambda_{\textsc{fpr}}$ and $\lambda_{\textsc{tpr}}$ obtain the following forms
$\lambda_{\textsc{fpr}}$ $=-\frac{1}{2s^{2}||x||^{2}}\langle x,x-m_{y}\rangle\left(\frac{2s}{\sqrt{S}}\langle x,z\rangle+\frac{1}{S}\langle x,x-m_{y}\rangle\right)$ (8)
$\lambda_{\textsc{tpr}}$ $\approx-\frac{1}{2s^{2}||x||^{2}}\langle x,x-m_{y}\rangle\left(\frac{2s}{\sqrt{S}}\langle x,z\rangle-\frac{2}{S}||x||^{2}+\frac{1}{S}\langle x,x-m_{y}\rangle\right),$ (9)
where $z\sim\mathcal{N}(0,I_{d})$.
Proof.
We defer the proof to Appendix C. ∎
The factors independent of $S$ in for both $\lambda_{\textsc{fpr}}$ and $\lambda_{\textsc{tpr}}$ correspond to the $-(\mu_{\textrm{in}}-\mu_{\textrm{out}})/(2\sigma^{2})$. Figure A.5 illustrates how this factor behaves as a function of $S$ in our experiments with the simple model. As predicted by the theory, the factor (i.e. the difference in means divided by the variance) remains fairly constant across different values of $S$.
Using Lemma 1 we can derive a rate for the decision threshold $\tau$:
Corollary 2.
In order to maintain constant fpr, we need to adjust $\tau$ as $O(1/\sqrt{S})$.
According to Corollary 2, the $\tau:=\tau(S)$ evolves as $O(1/\sqrt{S})$, and we will approximate it as $\tau(S)=\gamma_{1}\frac{1}{\sqrt{S}}$, with constant $\gamma_{1}$ depending on the target fpr.
Theorem 3.
$\log(\textsc{tpr}-\textsc{fpr})$ depends linearly on $\log(S)$.
Proof.
From Lemma 1, we see that $\lambda_{\textsc{fpr}}$ and $\lambda_{\textsc{tpr}}$ differ only in a single term which evolves as $O(1/S)$. Thus for sufficiently large $S$ we can consider this additional r.v. as a small constant shift to the distribution of $\lambda_{\textsc{fpr}}$. Let us denote this variable now with $q$. When $S$ grows, $q$ becomes small, and we have
[…]
$\approx p_{\lambda_{\textsc{fpr}}}(\tau)q.$ (11)
Taking the log yields
$\log(\textsc{tpr}-\textsc{fpr})$ $\approx\log\left(p_{\lambda_{\textsc{fpr}}}\left(\tau\right)\right)+\log q$ (12)
$=\log\left(p_{\lambda_{\textsc{fpr}}}\left(\tau\right)\right)-\log S+\gamma_{2}.$ (13)
From Lemma 1, we can see that a scaled r.v. $\hat{\lambda}_{\textsc{fpr}}:=\sqrt{S}\lambda_{\textsc{fpr}}$ becomes almost independent of $S$ for large enough $S$. Hence using the change of variables formula, we have that $p_{\lambda_{\textsc{fpr}}}(t)=p_{\hat{\lambda}_{\textsc{fpr}}}(\sqrt{S}t)\sqrt{S}$. Now, evaluating $p_{\lambda_{\textsc{fpr}}}(t)$ at $t=\tau(S)$, we get $p_{\lambda_{\textsc{fpr}}}(t)=p_{\hat{\lambda}_{\textsc{fpr}}}(\gamma_{1})\sqrt{S}$, where the $p_{\hat{\lambda}_{\textsc{fpr}}}(\gamma_{1})$ is independent of $S$. Therefore, ignoring additive constants,
$\log(\textsc{tpr}-\textsc{fpr})$ $\approx\log\sqrt{S}-\log S=\frac{1}{2}\log S-\log S=-\frac{1}{2}\log S.$ (14)
This shows that $\log(\textsc{tpr}-\textsc{fpr})$ depends linearly on $\log S$. ∎
Compared to the model in Section 3.3, our theoretical model predicts $\beta_{S}=-1/2$. Our empirical estimates match this for high fpr. For lower FPR, $\beta_{S}$ decreases, which indicates there may be finer details of $p_{\lambda_{\textsc{fpr}}}$ we are missing in this analysis.
5 MIA Vulnerability For Data Samples
This section extends our investigation from the data-set-level properties (Section 3) to individual data-sample-level properties. Previously Yu et al. (2023) have studied correlations between classes and privacy risk before, but in comparison to their study, we look at data samples independent from classes. We investigate two properties that could potentially influence the vulnerability of data points:
  • Per-sample gradients: Per-sample gradients form the basis for DP individual accounting (Feldman and Zrnic, 2021; Koskela et al., 2023). Under individual Gaussian DP (GDP, Dong et al., 2022) accounting (Koskela et al., 2023), the cumulative $\mu$-GDP privacy loss of individual $i$ with individual gradient norms $g_{i,t}$ is $\mu_{\text{i,tot}}\propto\sqrt{\sum_{t}g_{i,t}^{2}}$. Our hypothesis is that this similarly holds also for MIA vulnerability in non-DP settings.
  • Feature space distances: The pre-trained feature extractors project images into a feature space in which one can compute distances between individual data samples. Our hypothesis is that data samples that are further away from other data samples are easier to distinguish.
5.1 Experimental Setup
Our model training is done as in Section 3.1. The target models are fine-tuned pre-trained ViT-B (Head) and R-50 (Head) models. Below we describe how we measure the sample properties and individual tpr at fpr.
Recording of gradient norms During the training of the target models we record the per-sample gradient norm for every sample at every step (independent of being part of the current batch) efficiently using Opacus (Yousefpour et al., 2021). The recording of the gradient norms is identical to the DP individual accounting, with the exception that we record the gradient norms of out-samples as well. In our analysis, we are utilising the square root of the sum of squared gradient norm per sample over deciles of steps (i.e. the sum of squared individual gradient norms over the first 10% of steps, second 10% of steps, …, last 10% of steps).
$G_{i}^{(\alpha)}=\sqrt{\sum_{t=\lfloor T(\alpha-0.1)\rfloor}^{\lfloor T\alpha\rfloor}g_{i,t}^{2}},\;\alpha=(0.1,0.2,\ldots,1.0)$ (15)
Computation of distance in feature space All samples are projected into a feature space using the pre-trained model. We measure the Euclidean distance of all examples (in and out) to all in-samples in the feature space and report the distance to the nearest in-sample.
tpr at fpr for individual samples We study individual samples, but aggregate the results over multiple data samples by binning the samples based on their data sample properties to reduce the noise. For each data sample we compute its property and assign the sample to a bin based on its percentile in relation to other samples. We then compute the tpr at fpr for each bin. We finally report for each data property bin the percentile of the tpr at fpr.
Throughout our experiments we noticed that the amount of data points needs to be sufficiently large, in order to allow for an appropriate estimation of the tpr at fpr. Thus, we use a sufficient amount of shots ($S$) so that the size of in- and out-set combined is at least $10002|\mathcal{D}|=$ 1000 (see Table A10).
5.2 Individual Gradient Norms Are Predictive of Individual MIA Vulnerability
Figure 5 shows the observed relationships between the MIA vulnerability and the average individual gradient norms at $\textsc{fpr}=0.001$. When plotting the data for all deciles of steps at once, we find that the average gradient norms are correlated with the vulnerability in all our experiments, but the strength of the correlation depends on the data set. The $R^{2}$ (coefficient of determination) compares the observed relationship to a scenario where individual gradient norms would be a perfect predictor (corresponding to a scenario where all points are on the diagonal line). See more results in Table A10 and Figure A.6 in the Appendix.
Figure 5: MIA vulnerability (tpr at fpr; percentiles among a seed) as a function of individual gradient norms (binned among a seed). Data obtained over six seeds. The $R^{2}$ score compares the observed relationship to a scenario where individual gradient norms would be a perfect predictor.
Which parts of the training are more predictive?
In Figure 6, we examine whether the gradient norms at different stages of training are a better predictor of vulnerability by computing the ratio between the $R^{2}$ per iteration decile and $R^{2}$ over the whole training. For most data sets the ratio is constantly at a value of roughly 1, which suggests that gradient norms over even short segments can be as predictive as the norms of the whole trace.
Furthermore, there are deviations that indicate that for some data sets, gradient norms during the early iterations can be relatively more predictive than the norms of the entire trace. This is interesting, as the DP bound would directly depend on the norms of the entire trace.
Figure 6: Ratio between $R^{2}$ per iteration decile and $R^{2}$ over the whole training. A ratio $>1$ means that the decile is as more predictive than whole trace and a ratio $<1$ means that it is less predicitive.
5.3 Distance to Neighbors in Feature Space of Pre-trained Model
Figure 7 shows the observed relationships between the MIA vulnerability and distance to the nearest neighbor in the feature space. The distance of data samples to their nearest neighbor correlates with the vulnerability all analysed data sets but Patch Camelyon and Resics45. The observation holds over all fpr and for R-50 (Head) as well, refer to Figures A.9 and A.10 in the Appendix.
Figure 7: MIA vulnerability (tpr at fpr; percentiles among a seed) as a function of feature space distance (binned among a seed). The results have been obtained over six seeds.
@@ 6 Discussion
As the best means of protecting privacy, DP, reduces the utility of models, it is important to understand when it is necessary. Our results highlight that models are most vulnerable to MIA when the number of examples per class is low. This leads to an unfortunate dilemma: DP seems most important in the regime where it likely reduces the model utility the most.
@@
Our work measures the vulnerability through running one particular MIA. There is no guarantee that models that are not vulnerable to this attack might not be vulnerable to some stronger attack. Formal bounds on attack susceptibility would require something like DP.
Due to computational constraints, our predictive model is based only on transfer learning with fine-tuning only a part of the model. Limited comparison of our model predictions to previous results indicates it may underestimate their vulnerability.
Broader Impact
Our work systematically studies factors influencing privacy risk of trained ML models. This has significant positive impact on users training ML models on personal data by allowing them to understand and limit the risks.
Reproducibility
We provide the code for reproducing the experiments in the supplementary material and will publish the code in an open repository after acceptance of the paper. All used pre-trained models and data sets are publicly available.
Acknowledgements

### v3

@@ 4.2 Experimental results
Using the setting described above, we study how the number of classes and the number of shots affect the vulnerability (tpr at fpr as described in Section 2) using LiRA. We make the following observations:
  • A larger number of $S$ (shots) decrease the vulnerability in a power law relation as demonstrated in Figure 1(a). We provide further evidence of this in the Appendix (Figures A.2, A4 and A5).
  • Contrary, a larger number of $C$ (classes) increases the vulnerability as demonstrated in Figure 1(b) with further evidence in Figures A.3, A6 and A7 in the Appendix. However, the trend w.r.t. $C$ is not as clear as with $S$.
@@
[…]
RMIA In Figure 2 we compare the vulnerability of the models to LiRA and RMIA as a function of the number of $S$ (shots) at $\textsc{fpr}=0.1$. We observe the power-law for both attacks, but the RMIA is more unstable than LiRA (especially for lower fpr). More results for RMIA are in Figures A.6, A.7 and A.8 in the Appendix.
[…]
@@ 4.3 Model to predict dataset vulnerability
The trends seen in Figure 1 suggest the same power law relationship that we derived for the simplified model of membership inference in Section 3. We fit a linear regression model to predict $\log(\textsc{tpr}-\textsc{fpr})$ for each $\textsc{fpr}=10^{-k},k=1,\dots,5$ separately using the $\log C$ and $\log S$ as covariates with statsmodels (Seabold & Perktold, 2010). The general form of the model can be found in Equation 27, where $\beta_{S},\beta_{C}$ and $\beta_{0}$ are the learnable regression parameters.
@@
We utilise MIA results of ViT-B (Head) (see Table A4) as the training data. Based on the $R^{2}$ (coefficient of determination) score ($R^{2}=0.930$ for the model trained on $\textsc{fpr}=0.001$ data), our model fits the data extremely well. We provide further evidence for other fpr in Figure A.4 and Table A9 in the Appendix. Figure 3 shows the parameters of the prediction model fitted to the training data. For larger fpr, the coefficient $\beta_{S}$ is around $-0.5$, as our theoretical analysis predicts. However, the coefficient value decreases for small fpr. This is perhaps because the power-law in Equation (24) only holds for large $S$, and for small fpr Equation (24) significantly underestimates the vulnerability in small-$S$ regime (see Appendix D).
Figure 3: Coefficient values for different fpr when fitting a regression model based on Equation 27 fitted on data from ViT-B (Head) with LiRA (Table A4). The error bars display the 95% confidence intervals based on Student’s t-distribution.
@@
  • R-50 (FiLM): Figure 4(b) shows that the prediction quality is good for R-50 (FiLM) models. These models are fine-tuned with parameter-efficient FiLM (Perez et al., 2018) layers (See Section E.1). Tobaben et al. (2023) demonstrated that FiLM layers are a competitive alternative to training all parameters. We supplement the MIA results of Tobaben et al. (2023) with own FiLM training runs. Refer to Table A8 in the Appendix.
@@
(b) Evaluating prediction performance for target models 1 left: fine-tuned with FiLM (data partially from Tobaben et al. (2023), see Table A8) 2 right: trained from scratch (data from Carlini et al. (2022)) . The linear model is robust to changing the fine-tuning method from Head to FiLM, but from scratch training seems to be more vulnerable than predicted. Note that the results from Carlini et al. (2022) use data augmentation while we do not.
@@ 5 Discussion
As the best means of protecting privacy, differential privacy, reduces the utility of models, it is important to understand when it is necessary. Our results highlight that models are most vulnerable to MIA when the number of examples per class is low. This reinforces the well-known unfortunate dilemma: differential privacy seems most important in the regime where it likely reduces the model utility the most.
One major reason for MIA vulnerability can be memorisation of the training data. Feldman & Zhang (2020) experimentally test memorisation in neural network training, and find that according to their definition, a large fraction of the training data are memorised when training from scratch, while only few are when fine-tuning. This is aligned with our results that indicate from scratch training to be more vulnerable than fine-tuning.
@@
Our work measures the vulnerability through running two particular MIAs. There is no guarantee that models that are not vulnerable to this attack might not be vulnerable to some stronger attacks. Formal bounds on attack susceptibility would require something like differential privacy. Both our theoretical and empirical analysis focus on deep transfer learning using fine-tuning. Models trained from scratch are likely to be more vulnerable.
Acknowledgments